OpenStudy (anonymous):

Prove: arcsin[(x-1)/(x+1)] = 2arctan[sqrt(x)] - pi/2 Prove: arcsin[(x-1)/(x+1)] = 2arctan[sqrt(x)] - pi/2 @Mathematics

OpenStudy (anonymous):

\[\sin^{-1}[ (x-1)\div(x+1)] = 2\tan^{-1} (\sqrt{x})-\pi \div 2\]

OpenStudy (jamesj):

There are a few ways to skin this cat. My favorite today is this: Define a function \[ f(x) = \arcsin\left( \frac{x-1}{x+1} \right) - 2 \arctan\sqrt{x} \] Show that \( f'(x) = 0 \) and thus \( f(x) = constant \). Then show that constant is \(\pi/2\) (e.g., substituting \( x = 0 \) ).

OpenStudy (jamesj):

correction: constant = \( -\pi/2 \).

myininaya (myininaya):

what happens if we take sin( ) of both sides \[\sin^{-1}[ \frac{x-1}{x+1}] = 2\tan^{-1} (\sqrt{x})-\frac{\pi}{2}\] \[\sin(sin^{-1}[ \frac{x-1}{x+1}]) = sin(2\tan^{-1} (\sqrt{x})-\frac{\pi}{2})\] \[\frac{x-1}{x+1}=\sin(2 \tan^{-1}(\sqrt{x})-\frac{\pi}{2})\] \[\frac{x-1}{x+1}=\sin(2 \tan^{-1}(\sqrt{x})) \cos(\frac{\pi}{2})-\sin(\frac{\pi}{2})\cos(2 \tan^{-1}(\sqrt{x}))\] \[\frac{x-1}{x+1}=0-\cos(2 \tan^{-1}(\sqrt{x}))\] \[\frac{x-1}{x+1}=-\cos(2 \tan^{-1}(\sqrt{x}))\] assume we have right triangle now let \[u=\tan^{-1}(\sqrt{x}) => \tan(u)=\sqrt{x}=\frac{\sqrt{x}}{1} (=\frac{opp}{adj})\] \[hyp=\sqrt{x+1}\] so we have cos(2u) \[\cos(2u)=\cos^2(u)-\sin^2(u)=(\frac{1}{\sqrt{x+1}})^2-(\frac{\sqrt{x}}{\sqrt{x+1}})^2\] \[=\frac{1}{x+1}-\frac{x}{x+1}=\frac{1-x}{x+1}\] so we have \[\frac{x-1}{x+1}=-\frac{1-x}{x+1}\] which is true

OpenStudy (anonymous):

rofl no worries man tyvm :)

myininaya (myininaya):

woman*

myininaya (myininaya):

or lady

OpenStudy (anonymous):

oh my mistake

myininaya (myininaya):

or goddess will do

OpenStudy (anonymous):

lmao

myininaya (myininaya):

james's cal way sounds great too but i like trig its so neat-o

OpenStudy (anonymous):

kk ty

OpenStudy (anonymous):

i took the derivative like james said and i got to\[((x+1)^{2}-2\sqrt{x}) / (2x+2\sqrt{x})\] and i need to show that it is a constant

OpenStudy (anonymous):

u there goddess?

OpenStudy (anonymous):

sorry the bottom was \[2x \sqrt{x} + 2\sqrt{x}\]

OpenStudy (jamesj):

That's not quite right. The derivative of the arcsin term is \[ \frac{2}{(1+x)^2} \frac{1}{\sqrt{1 - \left(\frac{x-1}{x+1}\right)^2}} \] \[= \frac{1}{1+x}\frac{2}{\sqrt{(x+1)^2 - (x-1)^2 }} \] \[ = \frac{1}{1+x}\frac{2}{\sqrt{4x}} \] \[ = \frac{1}{\sqrt{x}(1+x)} \] Now I'll let you figure out the derivative of the arctan term again carefully and add it to this one.

OpenStudy (anonymous):

where did the 2/(x+1)^2 come from?

OpenStudy (jamesj):

that's the derivative of (x-1)/(x+1)

OpenStudy (anonymous):

oh ok thx

myininaya (myininaya):

let me know if you got it or not? i have part of the solution using jame's way you can compare if you want

myininaya (myininaya):

i guess i will go in post what i have \[\text{ let } u=\sin^{-1}(\frac{x-1}{x+1}) => \sin(u)=\frac{x-1}{x+1} (=\frac{opp}{hyp})\] =>\[adj=\sqrt{(x+1)^2-(x-1)^2}=\sqrt{x^2+2x+1-(x^2-2x+1)}=\sqrt{4x}=2 \sqrt{x}\] so anyways we are trying to find u' \[u' \cos(u)=\frac{1(x+1)-1(x-1)}{(x+1)^2}=\frac{x+1-x+1}{(x+1)^2}=\frac{2}{(x+1)^2}\] \[u'=\frac{2}{\cos(u) \cdot (x+1)^2}=\frac{2}{\frac{2 \sqrt{x}}{(x+1)}(x+1)^2}\] \[=\frac{x+1}{ \sqrt{x} (x+1)^2}\] ok now moving on to \[\text{ let } m=\tan^{-1}(\sqrt{x}) => \tan(m)=\sqrt{x}=\frac{\sqrt{x}}{1} (=\frac{opp}{adj})\] =>\[hyp=\sqrt{x+1}\] \[\sec^2(m) m'=\frac{1}{2 \sqrt{x}}\] \[m'=\frac{1}{2 \sqrt{x} \sec^2(m)} =\frac{1}{2 \sqrt{x} (\sqrt{x+1})^2}=\frac{1}{2 \sqrt{x}(x+1)}\] but we have 2m' \[2m'=\frac{1}{\sqrt{x}(x+1)}\] so f'(x)=\[\frac{x+1}{\sqrt{x}(x+1)^2}-\frac{1}{\sqrt{x} (x+1)}=\frac{x+1-(x+1)}{(x+1)^2 \sqrt{x}}=0\]

OpenStudy (anonymous):

ok looks good thank you both

OpenStudy (anonymous):

so for the arctan part i got \[1/(2\sqrt{x}(x+1))\] then you whould subtract them? \[1/(\sqrt{x}(x+1)) - 1/(2\sqrt{x}(x+1))\]

myininaya (myininaya):

2arctan part

OpenStudy (jamesj):

Look again at the function f. The coefficient of the arctan term is -2.

myininaya (myininaya):

yea what james said

OpenStudy (jamesj):

Hence \[ f'(x) = \frac{1}{\sqrt{x}(1+x)} - \frac{2}{2\sqrt{x}(1+x)} = 0\]

OpenStudy (anonymous):

ok now i get it :)